OpenStudy (anonymous):

Suppose that h(x)=(fog)(x)=f(g(x)) and that the following hold:? f(2) = 8, g(2)=2, f'(2)=1/2, g'(2)= -3, f(3)=3, g(3)=2, f'(3)=2 a) What is the derivative of f(g(x)) at x =2? b) If the derivative of f(g(x)) equals 3 at x=3, what is g'(3)?

OpenStudy (anonymous):

[f(g(x))]' = f'(g(x)) * g'(x) [f(g(2))]' = f'(g(2)) * g'(2) = f'(2) * g'(2) = (1/2) * (-3) = -3/2

OpenStudy (anonymous):

that's a...

OpenStudy (anonymous):

The question may be asking if it's f ' (g(x)) though.

OpenStudy (anonymous):

If that's the case, f ' (g(2)) = f ' (2) = 1/2

OpenStudy (anonymous):

f(g(3))]' = 3 = f'(g(3)) * g'(3) 3 = f'(2) * g'(3) 3 = (1/2) * g'(3) so g'(3) = 6

OpenStudy (anonymous):

Thanks! What rule did you incorporate in there?

OpenStudy (anonymous):

just the chain rule and only the chain rule...

OpenStudy (anonymous):

Oh, I had used the product rule in mine for some reason. Anyway, thank you so much!

OpenStudy (anonymous):

yw....:)